# Integrating Window-Based Correlated Decoding with Constant-Time Logical Gates for Large-Scale Quantum Computation

### v1

@@ Abstract
In contrast, the sliding windows works well in delayed fixup circuits. In a feed-forward sliding window architecture, the window with length $n_{w}$ slides forward in time (see Fig. 3a). At each step, correlated decoding is performed on $n_{w}$ rounds of syndrome data. The decoding result of each window is divided into two regions: a commit region and a buffer region with lengths $n_{c}$ and $n_{b}$, respectively, where $n_{w}=n_{c}+n_{b}$. After that, the window slides forward by $n_{c}$ rounds. In other words, the buffer region is the overlap between two adjacent windows. To ensure fault tolerance of the decoding, $n_{w}$, $n_{c}$, and $n_{b}$ are typically proportional to $\Theta(d)$.
@@
Since the syndrome checks near the window boundary lack future syndrome information, they might be matched incorrectly. Thus, only the decoding results within the commit region, which is sufficiently distant from the boundary, are considered reliable. Specifically, if all vertices of a hyperedge lie entirely within the commit region or spanning two regions, the result is preserved as final and the syndrome checks on the vertices are flipped. If all vertices of a hyperedge are entirely within the buffer region, it is deferred for the next window to resolve.
@@
Figure 3: Sliding window architecture and application in $T$-gadget. (a) Temporally feed-forward sliding window. Syndromes and hyperedges from each correlated decoding are represented within a window. The windows are processed sequentially, with each step removing the green hyperedge. (b) The measurement of |$\ket{T}$ occurs in the commit region of the second window, followed by the immediate measurement of |$\ket{S}$. In the third window, |$\ket{S}$ is independently decoded, with the state persisting for at most one window length.
@@
Then, we can discuss circuits containing mid-circuit measurements and feedback more concretely. Taking the $T$-gadget in Fig. 3b as an example, when state |$\ket{T}$ measurement appears in the commit region of a sliding window, the decoding yields a reliable measurement result. Following this, the state |$\ket{S}$ is measured in the appropriate basis immediately. The measurement outcome of |$\ket{S}$ will depend on the decoding result from the subsequent window. In particular, in the subsequent window, since the state |$\ket{S}$ does not interact with other qubits via two-qubit gates, it can be decoded independently as a single logical qubit. Note that independent decoding of |$\ket{S}$ states is crucial, especially when multiple non-commuting $T$ gates with $S$ fixup are performed consecutively. If the previous |$\ket{S}$ state is decoded within an entire window, the next |$\ket{S}$ measurement must be delayed until the following window. For a sequence of $S$ fixups, this will clearly result in an accumulation of errors.
@@
This example also highlights an additional advantage of window-based correlated decoding. When there are no two-qubit gate interactions between subsets of qubits within a window, the correlated decoding can be partitioned into several independent parts. This partitioning decreases the number of variables in the correlated decoding and enables further time reductions through parallelization. Essentially, this represents a specific instance of the spatial parallel window decoding, which will be discussed later.
Spatial Parallel Windows—Our starting point is the concept of TOQC for large-scale quantum computing, first proposed in Ref. [20]. Fig. 4a illustrates an circuit of TOQC, containing teleportation and delayed fixup circuits. The implementation time of TOQC is primarily dominated by the depth of measurements and classical feedback, thus reducing time overhead in large-scale computing tasks.
@@
Figure 4: Application of Spatial Windows in TOQC. (a) Parallel Clifford+$T$ gates in TOQC. The blue boxes divide the space into multiple blocks. The state |$\ket{S}$ in the $T$ gadget and its delayed measurement are also highlighted. (b) When the temporal window is shorter than the spacing between transversal CNOT gates, the window divides into independent parts naturally. (c) If the temporal window covers the entire circuit, it can be spatially divided, with syndrome data processed in a feed-forward manner. (d) In parallel spatial windows, decoding occurs in two stages, retaining green hyperedges as results in each.
@@
Here, similar to temporal windows, we find that TOQC circuits can be divided into spatial windows for correlated decoding. Notice that TOQC circuits can naturally be divided into blocks spatially, each comprising Bell state preparations, logical Clifford+$T$ gates, and Bell-basis measurements and feedback (see Fig. 4a). The blocks are interconnected by the initial and final transversal CNOT gates.
Now let us explain spatial windows in two scenarios (see Fig. 4b and 4c). First, if the depth $L$ of logical Clifford+$T$ gates in each block satisfies $L>n_{w}$ (recall that $n_{w}$ is the length of the temporal window), the initial and final transversal CNOT gates will not be in the same window. Consequently, this allows the temporal window to be divided into several independent parts for parallel correlated decoding. Conversely, if $L\leq n_{w}$, the temporal window cannot be directly divided into independent parts. In this case, we further divide the windows spatially, with each window containing syndrome data from $m_{w}$ blocks. Similarly, each window is composed of a commit region and a buffer region, containing $m_{c}$ and $m_{b}$ blocks, respectively. The intuition for spatial windows is that syndrome information between blocks that are spatially far apart becomes nearly independent. To keep the commit region away from the window boundary, the size of the buffer region is set to satisfy $m_{b}L=\Theta(d)$. In this setting, an error string connecting the commit region to the window boundary will contain at least $m_{b}L$ hyperedges. The condition $m_{b}L=\Theta(d)$ ensures such errors occur with exponentially low probability as $d$ increases and also confirms the comparable sizes of the temporal and spatial windows.
Spatial window decoding can be executed in a feed-forward manner. First, correlated decoding is applied to the syndrome data within the window, preserving hyperedges located entirely within the commit region or those spanning both the commit and buffer regions. Subsequently, the syndrome checks on the vertices of these hyperedges are flipped.
@@
Numerical Results—We present two examples (see details in Supplementary Material) to illustrate the performance of the window-based correlated decoding. In the first example, the length of temporal window is $n_{c}=n_{b}=\frac{d+1}{2}$. In the second, we test both feed-forward and parallel spatial windows with $m_{c}=1$ in both cases. Our setup ensures that the length of error strings from the commit region to the window boundary is at least $\frac{d+1}{2}$. Compared to correlated decoding on the syndrome data of the entire circuit, the performance of window-based correlated decoding is almost identical (see Fig. 5).
Figure 5: Decoding failure probabilities $p_{\text{fail}}$ with and without window architecture. The decoding methods are distinguished by colors, and the different code distances are represented by varying shades. (a) Comparison of $p_{\text{fail}}$ with and without temporal feed-forward window decoding. (b) Comparison of $p_{\text{fail}}$ with and without spatial window decoding, including both feed-forward and parallel types.
@@
Large-Scale Quantum Computation—Let us consider applying our architecture to large-scale quantum computation, using Shor’s algorithm as an example. Our analysis builds upon Ref. [21], comparing the time and space overhead of implementing Shor’s algorithm using fast transversal gates in ion trap systems versus the lattice surgery [33, 34, 22] in superconducting systems.
@@
Some basic assumptions include the following. First, it is assumed that the error rates for physical gates, initialization and measurements are comparable between the two systems, while the ion-trap system can provide required connectivity. Second, the QEC round time for the ion-trap system is 100 to 1,000 times that of the superconducting system [26, 27, 28]. Additionally, in ion trap systems, the reaction time, defined by measurement plus classical processing and feedback time, is about an order of magnitude longer than in superconducting systems due to longer measurement time [32, 35, 36, 37].
@@
Ref. [21] identifies the two most time-consuming subroutines as the ripple-carry adder [38] and the lookup table [39]. The ripple-carry adder, executable via TOQC circuits, is limited by the reaction time. At this point, fast transversal gates offer negligible time savings since the computation time is proportional to the measurement depth. In contrast, the implementation time of lookup table is dominated by multi-target CNOT gates, also known as Clifford-limited [32]. With fast transversal gates and ancilla cat states, multi-target CNOT gates can be implemented in a single QEC round, mitigating the effect of slower QEC rounds in ion trap systems by approximately a factor of $d$. Given the above factors, the total time of these two subroutines is approximately an order of magnitude longer in ion trap systems (see Supplementary Material for a more specific calculation).
@@
Moreover, fast transversal gates also reduce the space overhead. Since the algorithm is executed with fewer QEC rounds by fast transversal gates, the accumulated logical error rate on the computational qubits will be lower, thus relaxing the requirement for the code distance $d$. This reduction in space overhead is quite significant, approximately 14% when reducing the code distance $d$ from 27 to 25 (see Supplementary Material). Additionally, systems with flexible connectivity will further reduce qubit requirements by saving routing qubits and allowing flexible qubit reuse.
@@
However, the space occupied by magic state distillation must be considered [40]. The differing ratios of QEC round time and reaction time between ion trap and superconducting systems lead to variations in the generation and consumption rates of magic states. In ion trap system with fast transversal gates and TOQC circuits, magic states will be consumed faster, measured in units of QEC rounds, which indicates the distillation rate needs to increase. Fortunately, fast transversal gates applied in distillation can theoretically reduce space-time overhead by $\Theta(d)$ times [14]. Given that distillation protocol based on lattice surgery is highly optimized [41], the actual factor of reduction in space-time overhead will possibly be less than $d$. Considering these factors comprehensively, although distillation factories require more space, the overhead will remain acceptable. Overall, using fast transversal gates in ion trap systems yields space overhead of Shor’s algorithm comparable to using lattice surgery in superconducting systems [21], as corroborated by the calculations in the Supplementary Material.
Discussion—We present an architecture that integrates window-based correlated decoding with fast logical gates for universal quantum computing, offering broad applicability in large-scale quantum computing scenarios. The adaptation to temporal windows for parallelism, as seen in Refs. [18, 19], is straightforward. Nevertheless, investigating the general connection between spatial windows parallelism and the circuit structures remains an interesting question. Additionally, preliminary analysis of Shor’s algorithm in ion trap systems suggests promising potential, with more detailed resource evaluations and optimizations to be explored in future work.

### v2

@@ IV Correlated Decoding with Sliding Windows
In a feed-forward sliding window architecture, the window with length $n_{w}$ slides forward in time (see Fig. 4a). At each step, correlated decoding is performed on $n_{w}$ rounds of syndrome data. The decoding result of each window is divided into two regions: a commit region and a buffer region with lengths $n_{c}$ and $n_{b}$, respectively, where $n_{w}=n_{c}+n_{b}$. After that, the window slides forward by $n_{c}$ rounds. In other words, the buffer region is the overlap between two adjacent windows. To ensure fault tolerance of the decoding, $n_{w}$, $n_{c}$, and $n_{b}$ are typically proportional to $\Theta(d)$.
@@
Since the syndrome checks near the window boundary lack future syndrome information, they might be matched as a hyperedge incorrectly. Thus, only the decoding results within the commit region, which is sufficiently distant from the boundary, are considered reliable. Specifically, if all vertices of a hyperedge lie entirely within the commit region or spanning two regions, the result is preserved as final and the syndrome checks on the vertices are flipped. If all vertices of a hyperedge are entirely within the buffer region, it is deferred for the next window to resolve.
Next, let us describe in detail how correlated decoding is executed within a window. First, we need to establish the relationship between each error event $E_{j}$ with its error rate $p_{j}$ and its corresponding hyperedge $e_{j}$. For each error event $E_{j}$, a noiseless circuit with the error event $E_{j}$ is simulated, and the set of checks with a value of $-1$, denoted as $C(E_{j})$, is recorded. However, in a window of limited size, some checks in $C(E_{j})$ may fall outside the window. Thus, in correlated decoding within a specific window, we remove checks that lie outside the current window and consider the remaining checks as the vertices of hyperedge $e_{j}$. If none of the vertices are within the current window, the error event $E_{j}$ is removed from the error event set of the current window. Additionally, hyperedges spanning both the commit and buffer regions will be removed after current decoding. Therefore, for the error event $E_{j}$ and hyperedges $e_{j}$ that spans the commit and buffer regions of a window, we remove $E_{j}$ from the error event set in the next window, and only consider $e_{j}$ and $E_{j}$ in the current window.
After each correlated decoding within a window, we preserve the hyperedges inside the commit region and those on the boundary of the commit and buffer region as reliable results. The syndrome information corresponding to these hyperedges is then eliminated, meaning the check values at the vertices of these hyperedges are flipped. The decoding of subsequent windows is then carried out, and the final output consists of all reliable hyperedges in each decoding as the complete decoding result.
@@
Figure 4: Sliding window architecture and application in $T$-gadget. (a) Temporally feed-forward sliding window. The window slides from left to right along the temporal direction, with the spatial dimensions compressed into the other direction. The windows are processed sequentially, with each step removing the green hyperedges as reliable decoding results, while leaving the checks on the red hyperedges to be handled by the next window. (b) Application of the temporal window decoding in $T$-gadget. When the measurement of |$\ket{T}$ is located in the commit region of the second window, the decoder provides a reliable measurement outcome for |$\ket{T}$. Based on this outcome, the measurement basis is chosen, and the measurement of |$\ket{S}$ is subsequently performed. In the third window, |$\ket{S}$ is decoded independently. The |$\ket{S}$ state persists for at most one window length of syndrome extraction rounds.
@@
Then, we can discuss circuits containing mid-circuit measurements and feedback more concretely. Taking the $T$-gadget in Fig. 4b as an example, when state |$\ket{T}$ measurement appears in the commit region of a sliding window (the yellow one in the figure), the decoding yields a reliable measurement outcome. Following this, the state |$\ket{S}$ is measured in the appropriate basis immediately. The measurement outcome of |$\ket{S}$ will depend on the decoding result from the subsequent window. In particular, in the subsequent window, since the state |$\ket{S}$ does not interact with other qubits via two-qubit gates, it can be decoded independently as a single logical qubit. Note that independent decoding of |$\ket{S}$ states is crucial, especially when multiple non-commuting $T$ gates with $S$ fixup are performed consecutively. If the previous |$\ket{S}$ state is decoded within an entire window, the next |$\ket{S}$ measurement must be delayed until the following window. For a sequence of $S$ fixups, this will clearly result in an accumulation of errors.
@@
This example also highlights an additional advantage of window-based correlated decoding. When there are no two-qubit gate interactions between subsets of qubits within a window, the correlated decoding can be partitioned into several independent parts. This partitioning decreases the number of variables in the correlated decoding and enables further time reductions through parallelization. Essentially, this represents a specific instance of the spatial parallel window decoding, which will be discussed in the next section.
V Spatial Parallel Windows
@@
Figure 5: Application of Spatial Windows in TOQC. (a) Parallel Clifford+$T$ gates in TOQC. The blue boxes divide the space into multiple blocks. The state |$\ket{S}$ in the $T$ gadget and its delayed measurement are also highlighted. (b) When the temporal window is shorter than the spacing between transversal CNOT gates, the window divides into independent parts naturally (by the green dot lines in the figure). (c) If the temporal window covers the entire circuit, it can be spatially divided, with syndrome data processed in a feed-forward manner. (d) In parallel spatial windows, decoding occurs in two stages, retaining green hyperedges as results in each.
Our starting point is the concept of TOQC for large-scale quantum computing, first proposed in Ref. [24]. Fig. 5a illustrates an example of the TOQC circuit. Using teleportation and delayed fixup circuits, multiple groups of Clifford+$T$ gates can be executed in parallel, at the cost of requiring additional qubits. The implementation time of TOQC is primarily dominated by the depth of measurements and classical feedback, thus reducing time overhead in large-scale computing tasks. For more information on TOQC, please refer to the Appendix A.
@@
Similar to temporal windows, we find that TOQC circuits can be divided into spatial windows for correlated decoding. Notice that TOQC circuits can naturally be divided into blocks spatially, each comprising Bell state preparations, logical Clifford+$T$ gates, and Bell-basis measurements and feedback (see Fig. 5a). The blocks are interconnected by the initial and final transversal CNOT gates.
Now let us explain spatial windows in two scenarios. First, if the depth $L$ of logical Clifford+$T$ gates in each block satisfies $L>n_{w}$ (recall that $n_{w}$ is the length of the temporal window), the initial and final transversal CNOT gates will not be in the same window (see Fig. 5b). Consequently, this allows the temporal window to be divided into several independent parts for parallel correlated decoding. Conversely, if $L\leq n_{w}$, the temporal window cannot be directly divided into independent parts (see Fig. 5c). In this case, we further divide the windows spatially, with each window containing syndrome data from $m_{w}$ blocks. Similarly, each window is composed of a commit region and a buffer region, containing $m_{c}$ and $m_{b}$ blocks, respectively. The intuition for spatial windows is that syndrome information between blocks that are spatially far apart becomes nearly independent. To keep the commit region away from the window boundary, the size of the buffer region is set to satisfy $m_{b}L=\Theta(d)$. In this setting, an error string connecting the commit region to the window boundary will contain at least $m_{b}L$ hyperedges. The condition $m_{b}L=\Theta(d)$ ensures such errors occur with exponentially low probability as $d$ increases.
Spatial window decoding can be executed in a feed-forward manner. First, correlated decoding is applied to the syndrome data within the window, preserving hyperedges located entirely within the commit region or those spanning both the commit and buffer regions as the reliable decoding results. Then the syndrome checks on the vertices of these hyperedges are flipped. Subsequently, decoding is performed for the next window in a similar manner.
@@
Here, it is beneficial to discuss the size of both temporal and spatial windows. For temporal windows without TOQC implementation, their size is immediately apparent. Each temporal window processes syndrome data from $\Theta(d)$ rounds on $n$ qubits at a time. We assume that such size of correlated decoding remains acceptable. However, in TOQC circuits, a spatial window includes multiple $n$-qubit blocks. One might wonder whether the size of a spatial window still remains acceptable.
To ensure fault tolerance, the distance between the commit region and the window boundary must be sufficiently large. To be precise, the distance refers to the minimum number of hyperedges required to connect two vertices. In fact, this distance only depends on the size of the buffer region. For temporal and spatial windows, $n_{b}$ and $m_{b}$ can be set to $(d+1)/2$ and $(d+1)/2L$, respectively. This ensures that the probability of an error chain connecting the commit region to the window boundary is $\mathcal{O}(p^{(d+1)/2})$, which will not significantly contribute to the logical error rate $P_{L}$. On the other hand, the size of the commit region is not restricted. Typically, to output as many reliable decoding results as possible while keeping the overall window size manageable, the size of the commit region is set equal to that of the buffer region.
Finally, we would like to emphasize that spatial windows does not apply to more general circuits. Our work only demonstrates its applicability specifically to the structure of TOQC circuits. This is because, in TOQC circuits, the CNOT gates between blocks are very limited and appear only at the beginning and end of the circuit, which means that the correlation of error syndromes is influenced by spatial distance.
VI Numerical Results
Figure 6: Numerical simulation of two examples. (a-d) show the circuits we simulated along with the corresponding window-based decoding methods. (e-h) present the results of the numerical simulations. The statistical error of one standard deviation for each data point in the figure is no greater than the size of the marker for that point. (a) Circuit and temporal feed-forward sliding window for the first example. Each window contains $d+1$ rounds of syndrome data. (b) Circuit for the second example, with the similar structure as the TOQC. (c) Feed-forward spatial windows for decoding the second example. Each spatial window contains syndrome data on two logical qubits. (d) Parallel spatial windows for decoding the second example. (e) Decoding failure rate $p_{\text{fail}}$ at different code distances as a function of the physical error rate without window-based decoding for the first example. (f) The ratio of decoding failure rates with temporal windows to those without. (g) Decoding failure rate $p_{\text{fail}}$ at different code distances as a function of the physical error rate without window-based decoding for the second example. (h) The ratio of decoding failure rates with spatial windows to those without, including both feed-forward and parallel types.
We present two examples (see Fig. 6) to illustrate the performance of the window-based correlated decoding. Here, the logical qubits in the circuit are encoded on surface codes, enabling transversal CNOT gates and single-qubit Clifford gates from the set $\{X,Y,Z,H\}$. Fig. 6 (a-b) illustrate the specific circuits for both examples, where syndrome extraction is performed for single round after each logical gate. The division of windows is shown in Fig. 6a and (c-d). We also provide details of the noise model and simulation methods in Appendix  B.
In the second example, the circuit consists of five logical qubits and has a structure similar to TOQC with $L=(d+1)/2$. We tested the decoding process under two window architectures, feed-forward and parallel windows. In the feed-forward window architecture, each window contains syndrome data from two blocks. Correlated decoding is performed sequentially across four windows, with the window sliding by one block in the spatial direction each time. In the parallel window architecture, decoding is carried out in two steps: first, two disjoint spatial windows are decoded, followed by independent correlated decoding in the remaining buffer regions of the two parts.
In both two examples, our setup ensures that the distance from the commit region to the window boundary is at least $\frac{d+1}{2}$. This condition ensures that the additional logical error rate introduced by the windowing method is $\mathcal{O}(p^{(d+1)/2})$, on the same order as the logical error rate of the surface code encoding itself. For comparison to window-based decoding, we also performed correlated decoding without windows, meaning a single decoding process is applied to all syndrome data from the entire circuit. We define the decoding failure rate $p_{\text{fail}}$ as the probability of a logical Pauli error occurring on any logical qubit after decoding. Fig. 6e and Fig. 6g show the failure rates of decoding without the window architecture for the two examples. In comparison, Fig. 6f and Fig. 6h display the ratio of the failure rates with and without the window architecture for the same examples. The results indicate that the performance of window-based correlated decoding is almost identical, with only a very slight increase in $p_{\text{fail}}$.
VII Application in Large-Scale Quantum Computation
@@
In this section, we aim to provide a preliminary analysis of the impact of constant-time quantum gates on large-scale quantum computation, using the Shor algorithm as an example. Our analysis builds upon Ref. [25], comparing the time and space overhead of implementing Shor’s algorithm using constant-time transversal gates in ion trap systems versus the lattice surgery [34, 35, 26] in superconducting systems.
@@
Some basic assumptions include the following. First, it is assumed that the error rates for physical gates, initialization and measurements are equal for the two systems, while the ion-trap system can provide the required connectivity. Second, the QEC round time for the ion-trap system is 100 to 1,000 times that of the superconducting system [29, 15, 30]. Additionally, in ion trap systems, the reaction time, defined by measurement plus classical processing and feedback time, is about an order of magnitude longer than in superconducting systems due to longer measurement time [38, 41, 42, 43].
@@
Ref. [25] identifies the two most time-consuming subroutines as the ripple-carry adder [44] and the lookup table [45], both of which are also important subroutines in other algorithm [46, 47]. The ripple-carry adder, executable via TOQC circuits, is limited by the reaction time. At this point, fast transversal gates offer negligible time savings since the computation time is proportional to the measurement depth. In contrast, the implementation time of lookup table is dominated by multi-target CNOT gates, also known as Clifford-limited [38]. With fast transversal gates and ancilla cat states, multi-target CNOT gates can be implemented in a single QEC round, mitigating the effect of slower QEC rounds in ion trap systems by approximately a factor of code distance $d$. Given the above factors, the total time of these two subroutines is approximately an order of magnitude longer in ion trap systems (see Appendix D for a more specific calculation).
@@
Moreover, fast transversal gates also reduce the space overhead. Since the algorithm is executed with fewer QEC rounds by fast transversal gates, the accumulated logical error rate on the computational qubits will be lower, thus relaxing the requirement for the code distance $d$. This reduction in space overhead is approximately 14% when reducing the code distance $d$ from 27 to 25 (see Appendix D). Additionally, systems with flexible connectivity will further reduce qubit requirements by saving routing qubits and allowing flexible qubit reuse.
@@
However, the space occupied by magic state distillation must be considered [48]. The differing ratios of QEC round time and reaction time between ion trap and superconducting systems lead to variations in the generation and consumption rates of magic states. In ion trap system with fast transversal gates and TOQC circuits, magic states will be consumed faster, measured in units of QEC rounds, which indicates the distillation rate needs to increase. Fortunately, fast transversal gates applied in distillation can theoretically reduce space-time overhead by $\Theta(d)$ times [18]. Given that distillation protocol based on lattice surgery is highly optimized [49], the actual factor of reduction in space-time overhead will possibly be less than $d$. Considering these factors comprehensively, although distillation factories require more space, the overhead will remain acceptable. Overall, using fast transversal gates in ion trap systems yields space overhead of Shor’s algorithm comparable to using lattice surgery in superconducting systems [25], as corroborated by the calculations in Appendix D.
VIII Conclusion and Discussion
In this work, we present an architecture that integrates window-based correlated decoding with constant-time logical gates for universal quantum computing. Compared to the original correlated decoding, window-based correlated decoding significantly enhances scalability and efficiency, offering broad applicability for constant-time logical gates in large-scale quantum computing scenarios.
When discussing time-based windows, we only considered feedforward windows and do not address parallel windows. However, we also note that such parallelism for temporal windows has already been achieved in previous works on window-based decoding, as seen in Refs. [22, 23]. For window-based correlated decoding, there are no fundamental difficulties in applying this approach. Hence, adapting temporal windows for parallelism in our work remains feasible. On the contrary, as we mentioned, applying spatial windows and achieving parallelism is not always feasible for general circuit structures. Intuitively, the division of spatial windows appears to be related to the degree of transverse CNOT connectivity between the blocks of the circuit. Investigating the connection between spatial windows parallelism and the circuit structures remains an interesting question.
Additionally, preliminary analysis of Shor’s algorithm in ion trap systems suggests promising potential, indicating the value of constant-time logical gates in FTQC. However, we also acknowledge that a particularly detailed and accurate analysis is beyond the scope of this work. Therfore, we leave more detailed resource evaluations and optimizations to be explored in future work.